# Kernel vs. Kernel: Exploring How the Data Structure Affects Neural Collapse

### v2

@@ Abstract
Recently, a vast amount of literature has focused on the “Neural Collapse” (NC) phenomenon, which emerges when training neural network (NN) classifiers beyond the zero training error point. The core component of NC is the decrease in the within-class variability of the network’s deepest features, dubbed as NC1. The theoretical works that study NC are typically based on simplified unconstrained features models (UFMs) that mask any effect of the data on the extent of collapse. In this paper, we provide a kernel-based analysis that does not suffer from this limitation. First, given a kernel function, we establish expressions for the traces of the within- and between-class covariance matrices of the samples’ features (and consequently an NC1 metric). Then, we turn to focus on kernels associated with shallow NNs. First, we consider the NN Gaussian Process kernel (NNGP), associated with the network at initialization, and the complement Neural Tangent Kernel (NTK), associated with its training in the “lazy regime”. Interestingly, we show that the NTK does not represent more collapsed features than the NNGP for prototypical data models. As NC emerges from training, we then consider an alternative to NTK: the recently proposed adaptive kernel, which generalizes NNGP to model the feature mapping learned from the training data. Contrasting our NC1 analysis for these two kernels enables gaining insights into the effect of data distribution on the extent of collapse, which are empirically aligned with the behavior observed with practical training of NNs.
@@ 1 Introduction
Deep Neural Network classifiers are often trained beyond the zero training error point [1, 2]. In this regime, a phenomenon dubbed “Neural Collapse” (NC) emerges [3]. NC is typically described by the following components: (NC1) the networks’ deepest features exhibit a significant decrease in the variability of within-class samples, (NC2) the mean features of different classes approach a certain symmetric structure, and (NC3) the last layer’s weights become more aligned with the penultimate layer features’ means. This behavior has been observed both when using the cross-entropy (CE) loss [3] and the mean squared error (MSE) loss [4].
@@
Recently, a vast amount of literature has been dedicated to exploring NC (as surveyed in [5]), studying the effect of imbalanced data [6, 7], depthwise evolution [8, 9, 10, 11], fine-grained structures [12, 13, 14], and implications [15, 16, 17]. Note that without a sufficient decrease in the features’ within-class variability around the class means, measured by NC1 metrics, one may not gain valuable insights from the structure of the means of different classes. Therefore, oftentimes, NC papers focus specifically on NC1 rather than on other components of NC [12, 13, 14, 16, 11, 18]. Notably, most of the works that attempt to theoretically analyze the NC behavior [19, 6, 15, 4, 20, 8, 21, 7, 22, 12, 23, 10, 24] are based on variants of the unconstrained features model (UFM) [19], which treats the deepest features of the training samples as free optimization variables. Therefore, these analyses cannot predict the effect of the data structure on the extent of collapse.
@@
Since theoretically analyzing the behavior of (deep) NNs is challenging, simplifying approaches that are based on kernel methods [25] have gained massive popularity [26, 27, 28, 29, 30, 31]. Prominent examples include the NN Gaussian Process kernel (NNGP) [26, 27, 32], associated with the infinitely wide NNs at initialization, and the Neural Tangent Kernel (NTK) [28], associated with their training in the “lazy regime”, where the learning rate is sufficiently small. These approaches, and in particular NTK, were used to provide mathematical reasoning for deep learning phenomena such as achieving zero training loss [28, 30, 31], faster learning of lower frequencies [33], benefits of ResNets over fully connected networks [34, 35, 36], usefulness of positional encoding in coordinated-based NNs [37], and more. More recently, finite width variants of the aforementioned kernels have been studied [38, 39, 40], aiming to mitigate the gap between practical NNs behavior and infinite width analyses [41, 42, 43, 44, 45].
In this paper, we provide a kernel-based analysis for NC1, the core component of NC, which does not suffer from the limitations of UFM-based analysis. Thus, it allows us to explore how the data structure affects the collapse. Since kernels provide fixed feature mapping, we propose a “kernel vs. kernel” analysis — that is, gaining insights by comparing the properties across NN-related kernels.
Our main contributions can be summarized as follows:
@@
  • Given an arbitrary kernel function, we establish expressions for the traces of the within- and between-class covariance matrices of the samples’ features — and consequently, an NC1 metric that depends on the features only through the kernel function.
@@
  • We specialize our kernel-based NC1 to kernels associated with shallow NNs. We analyze it for NNGP (NN at initialization), and NTK (NN trained in the “lazy regime”) and show that, perhaps surprisingly, the NTK does not represent more collapsed features than the NNGP for prototypical data models.
  • As NC emerges from training, we consider an alternative to NTK: the recently proposed adaptive kernel [40], which generalizes NNGP to model the feature mapping learned from the training data. Contrasting our NC1 analysis for these two kernels enables gaining deeper insights into the effect of data distribution on the extent of collapse, which are empirically aligned with the behavior observed with practical training of NNs.
@@ 2 Related Work
As mentioned in the previous section, most of the works that attempt to analyze the NC behavior theoretically [19, 6, 15, 4, 20, 8, 21, 7, 22, 23, 10] are based on variants of the UFM [19]. The work in [12] attempts to generalize the UFM by adding a penalty term to the loss that ensures that the features matrix is in the vicinity of a predefined matrix. Yet, the model still lacks an explicit connection between the features and the data. In [46], the authors avoid optimizing the features directly but assume that the model is linear and that the data is nearly orthogonal, which are restrictive assumptions. In [47], the authors claim that having an exact class-wise block structure in the Gram matrix of the empirical NTK on training samples implies NC. Yet, they do not provide reasoning for reaching this collapsed Gram matrix and the analysis is still disconnected from the data. Here, however, we fully depart from the UFM approach and provide analysis that explicitly depends on the data. Furthermore, unlike most works, our analysis is applicable to the less studied case where the data is class imbalanced [6, 17, 23, 7]. Our kernel-based analysis utilizes results on NNGP and NTK from [48, 49, 27, 28, 31]. To simplify the analysis, we theoretically analyze kernels associated with shallow fully connected NNs. Focusing on shallow networks is justified by recent works demonstrating monotonic depthwise evolution of NC1 [8, 9, 10, 11, 12]. Specifically, a data structure that promotes a larger reduction in NC1 for shallow NNs is expected to be more collapsed when using deep NNs. Since NC is related to training, and we show the limitation of NTK to capture it when compared to NNGP, we also utilize the generalization of NNGP that has been proposed in [40]. In this adaptive kernel model, there is an explicit kernel function that depends on the training data. Recently, this richer model has been used to study phase transition behaviors, such as grokking [50], that cannot be captured with the data-independent kernels.
@@ 3 Problem Setup
$\bullet$ Neural Network: Unless stated otherwise, we consider a 2-layer fully connected neural network (2L-FCN) $\psi:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{2}}$ with $l^{th}$ layer width $d_{l},l\in\{1,2\}$, and point-wise activation function $\phi(\cdot):\mathbb{R}\to\mathbb{R}$. Let $\mathbf{W}^{(l)}\in\mathbb{R}^{d_{l}\times d_{l-1}},\mathbf{b}^{(l)}\in\mathbb{R}^{d_{l}}$ denote the weight and bias parameters of the $l^{th}$ layer. At initialization, the entries $W_{ij}^{(l)},b_{i}^{(l)}$ are drawn i.i.d from Gaussian distributions of mean 00 0 and variance $\sigma_{w}^{2}/d_{l-1},\sigma_{b}^{2}$, respectively. For an input $\mathbf{x}\in\mathbb{R}^{d_{0}}$ to the network $\psi(\cdot)$, we denote the $i^{th}$ component of the output vector $\hat{y}_{i}(\mathbf{x})\in\mathbb{R}$ as follows:
@@
$\displaystyle\hat{\mathbf{Y}}=\begin{bmatrix}\psi(\mathbf{x}^{1,1})&\cdots&\psi(\mathbf{x}^{C,n_{C}})\end{bmatrix}=\begin{bmatrix}\hat{\mathbf{y}}^{1,1}&\cdots&\hat{\mathbf{y}}^{C,n_{C}}\end{bmatrix},\hskip 20.0pt\mathbf{Y}=\begin{bmatrix}\mathbf{y}^{1,1}&\cdots&\mathbf{y}^{C,n_{C}}\end{bmatrix}$ (2)
@@
Note that training deep classifiers with MSE loss has been shown to be a useful strategy [51, 4], which is commonly considered in NC analyses [4, 8, 21].
@@ 4 Within-Class Variability Metric (NC1) for Kernels
In this section, we derive an NC1 metric that depends on the features only through the kernel function. Let $\mathbf{H}$ be a matrix encapsulating arbitrary feature vectors associated with samples of the $C$ classes. We define the within-class covariance $\boldsymbol{\Sigma}_{W}(\mathbf{H})$ and between-class covariance $\boldsymbol{\Sigma}_{B}(\mathbf{H})$ matrices of the features as follows:
@@
Based on these formulations, we define the variability metric $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$, introduced in [12] and used also in [14, 46, 52], as:
@@
5 Activation Variability in the Lazy Learning Regime
@@ 5.1 Limiting NNGP Kernel
Under the NN model and initialization stated in Section 3, as the hidden layer width $d_{1}\to\infty$, we can characterize the pre-activation kernel $K^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})$ in terms of the GP limit [26] (commonly referred to as the NNGP limit [27]) as follows:
@@
In this limit, the post-activation kernel $Q^{(1)}_{GP}(\cdot,\cdot)$ can have a closed form representation depending on the choice of activation function $\phi(\cdot)$ [27, 31]. The expression for Erf activation [48] is:
@@
The formulation for ReLU-based kernel $Q_{GP-ReLU}^{(1)}(\cdot,\cdot)$ [49] is presented in the Appendix B. Given a kernel function $Q(\cdot,\cdot)$ and samples $\mathbf{X}$, we can formulate the kernel Gram matrix $\mathbf{Q}\in\mathbb{R}^{N\times N}$ as:
@@ 5.2 Limiting NTK
In the infinite width limit, we also analyze the NTK [28] to understand the effect of optimization on the NN’s features in the “lazy regime”. Specifically, a well-known result is that in the infinite width limits (with initialization as per Section 3), the deepest feature mapping of the NN is fixed during gradient descent optimization with small enough learning rate, and is characterized by the NTK.
Formally, the recursive relationship between the NTK and NNGP [28, 31] can be given as follows:
@@
Here, $K_{GP}^{(2)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})$ can be defined using the recursive formulation [27, 28]:
@@
Remark on the limiting NTK. Denote by $w,\phi$ the parameters and the activation function of an $L$-layer NN $\psi(\cdot)$. The limiting NNGP is defined directly on the product of neurons, $\mathbb{E}_{w}\langle\phi(z_{i}(\mathbf{x})),\phi(z_{i}(\tilde{\mathbf{x}}))\rangle$, and provides kernel expressions for the inner product of features at different layers of the network (via recursion similar to (11)). On the other hand, the limiting NTK is defined as the inner product of the output’s gradients $\mathbb{E}_{w}\langle\nabla_{w}\psi(\mathbf{x}),\nabla_{w}\psi(\tilde{\mathbf{x}})\rangle$. The NTK theory shows that this can model the inner product only of the deepest features (i.e the output of the penultimate layer) [^1] Note that we denote NNGP $(Q_{GP}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j}))$ and NTK ($\Theta^{(2)}_{NTK}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})$) with different superscripts to be consistent with the literature. Yet both are associated with the output of the single hidden layer. .
5.3 1-D Gaussian Data with 2 Classes

### v3

# Can Kernel Methods Explain How the Data Affects Neural Collapse?

@@ Abstract
A vast amount of literature has recently focused on the “Neural Collapse” (NC) phenomenon, which emerges when training neural network (NN) classifiers beyond the zero training error point. The core component of NC is the decrease in the within-class variability of the network’s deepest features, dubbed as NC1. The theoretical works that study NC are typically based on simplified unconstrained features models (UFMs) that mask any effect of the data on the extent of collapse. To address this limitation of UFMs, this paper explores the possibility of analyzing NC1 using kernels associated with shallow NNs. We begin by formulating an NC1 metric as a function of the kernel. Then, we specialize it to the NN Gaussian Process kernel (NNGP) and the Neural Tangent Kernel (NTK), associated with wide networks at initialization and during gradient-based training with a small learning rate, respectively. As a key result, we show that the NTK does not represent more collapsed features than the NNGP for Gaussian data of arbitrary dimensions. This showcases the limitations of data-independent kernels such as NTK in approximating the NC behavior of NNs. As an alternative to NTK, we then empirically explore a recently proposed data-aware Gaussian Process kernel, which generalizes NNGP to model feature learning. We show that this kernel yields lower NC1 than NNGP but may not follow the trends of the shallow NN. Our study demonstrates that adaptivity to data may allow kernel-based analysis of NC, though, further advancements in this area are still needed. A nice byproduct of our study is showing both theoretically and empirically that the choice of nonlinear activation function affects NC1 (with ERF yielding lower values than ReLU). The code is available at: https://github.com/kvignesh1420/shallow_nc1.
@@ 1 Introduction
Deep Neural Network classifiers are often trained beyond the zero training error point (Hoffer et al., 2017; Ma et al., 2018). In this regime, a phenomenon dubbed “Neural Collapse” (NC) emerges (Papyan et al., 2020). NC is typically described by the following components: (NC1) the networks’ deepest features exhibit a significant decrease in the variability of within-class samples, (NC2) the mean features of different classes approach a certain symmetric structure, and (NC3) the last layer’s weights become more aligned with the penultimate layer features’ means. This behavior has been observed both when using the cross-entropy (CE) loss (Papyan et al., 2020) and the mean squared error (MSE) loss (Han et al., 2022).
@@
Recently, a vast amount of literature has been dedicated to exploring NC (as surveyed in (Kothapalli, 2023)), studying the effect of imbalanced data (Fang et al., 2021; Thrampoulidis et al., 2022), depthwise evolution (Tirer & Bruna, 2022; Rangamani et al., 2023; Sukenik et al., 2023; He & Su, 2023), fine-grained structures (Tirer et al., 2023; Yang et al., 2023; Kothapalli et al., 2023), and implications (Zhu et al., 2021; Galanti et al., 2022; Yang et al., 2022). Note that without a sufficient decrease in the features’ within-class variability around the class means, measured by NC1 metrics, one may not gain valuable insights from the structure of the means of different classes. Therefore, oftentimes, NC papers focus specifically on NC1 rather than on other components of NC (Tirer et al., 2023; Yang et al., 2023; Kothapalli et al., 2023; Galanti et al., 2022; He & Su, 2023; Xu & Liu, 2023). Notably, most of the works that attempt to theoretically analyze the NC behavior (Mixon et al., 2020; Fang et al., 2021; Zhu et al., 2021; Han et al., 2022; Ji et al., 2022; Tirer & Bruna, 2022; Zhou et al., 2022; Thrampoulidis et al., 2022; Yaras et al., 2022; Tirer et al., 2023; Dang et al., 2023; Sukenik et al., 2023; Wojtowytsch et al., 2020) are based on variants of the unconstrained features model (UFM) (Mixon et al., 2020), which treats the deepest features of the training samples as free optimization variables. Therefore, such UFM analyses rely solely on the labels and cannot predict the effect of training data on the extent of collapse.
@@
Since theoretically analyzing the behavior of (deep) NNs is challenging, simplifying approaches that are based on kernel methods (Schölkopf et al., 2002) have gained massive popularity (Neal, 1995; Lee et al., 2018; Jacot et al., 2018; Chizat et al., 2019; Arora et al., 2019; Lee et al., 2019). Prominent examples include the NN Gaussian Process kernel (NNGP) (Neal, 1995; Lee et al., 2018; Matthews et al., 2018), associated with the infinitely wide NNs at initialization, and the Neural Tangent Kernel (NTK) (Jacot et al., 2018), associated with their training in the “lazy regime”, where the learning rate is sufficiently small. These approaches, and in particular NTK, were used to provide mathematical reasoning for deep learning phenomena such as achieving zero training loss (Jacot et al., 2018; Arora et al., 2019; Lee et al., 2019), faster learning of lower frequencies (Ronen et al., 2019), benefits of ResNets over fully connected networks (Huang et al., 2020; Tirer et al., 2022; Barzilai et al., 2022), usefulness of positional encoding in coordinated-based NNs (Tancik et al., 2020), and more. More recently, finite width variants of the aforementioned kernels have been studied (Hanin & Nica, 2019; Lee et al., 2020; Seroussi et al., 2023), aiming to mitigate the gap between practical NNs behavior and infinite width analyses (Woodworth et al., 2020; Ghorbani et al., 2020; Wei et al., 2019; Li et al., 2020).
Why Kernels on Gaussian data for studying NC?
In this paper, we provide a kernel-based analysis for NC1 (the core component of NC) to bypass the limitations of UFM-based analysis. More importantly, our work aims to highlight the role of data in the study of NC and thus leverages the Gaussian datasets for a systematic theoretical and empirical analysis of NC1 with kernels. This setup allows us to explore how different aspects of the data such as dimension, sample size, class imbalance, and linear separability affect the collapse in a controlled fashion. To this end, as also mentioned above, we focus on (1) data-independent kernels such as NNGP/NTK whose functional form does not change with data, and (2) data-aware GP kernels (Seroussi et al., 2023) which model the feature mappings learned by finite-width shallow NNs (especially shallow Fully Connected Networks (FCNs)). Our main contributions can be summarized as follows:
@@
  • Given an arbitrary kernel function, we establish expressions for the traces of the within- and between-class covariance matrices of the features — and consequently, an NC1 metric that depends on the features only through the kernel function.
@@
  • We specialize our kernel-based NC1 to kernels associated with shallow NNs. We analyze it for NNGP and NTK and show that, perhaps surprisingly, the NTK does not represent more collapsed features than the NNGP for Gaussian data.
  • We analyze the data-aware GP kernel (Seroussi et al., 2023), which generalizes NNGP to model the feature mapping learned from the training data by shallow NNs. We present empirical results on Gaussian data to show that the adaptivity of such kernels yields lower NC1 and allows them to reasonably approximate the NC1 behavior of shallow NNs for linearly separable datasets. However, they may fail in regimes with imbalanced and non-linearly separable data in higher dimensions.
  • A byproduct of our study is showing both theoretically and empirically that the choice of the nonlinear activation function affects NC1. Specifically, we show that using ERF yields lower NC1 (more collapsed features) than using ReLU.
Paper Organization.
We review previous works on the UFM and kernels in the context of NC in Section 2. Section 3 introduces the setup based on 2 layer FCNs and kernels used throughout the paper. Section 4 introduces a generic NC1 metric for any arbitrary kernel function. This formulation is then used in Section 5 to theoretically and empirically analyze the NC1 of NNGP and NTK kernels. Note that the NTK feature map considers the FCN architecture and the GD optimization process (although not data-aware) for the NC1 analysis, and allows us to gain insights beyond NNGP (Seleznova et al., 2023). Owing to the limitations of the NTK, we wish to contrast NNGP with a kernel that takes into account both optimization and data. To this end, we transition to the recently introduced data-aware GP kernels by (Seroussi et al., 2023) for FCNs in Section 6. Finally, we discuss the limitations of such kernels in Section 7 and the conclusion in Section 8.
@@ 2 Related Work
The majority of the works that attempt to analyze NC behavior theoretically (Mixon et al., 2020; Fang et al., 2021; Zhu et al., 2021; Han et al., 2022; Ji et al., 2022; Tirer & Bruna, 2022; Zhou et al., 2022; Yaras et al., 2022; Dang et al., 2023; Sukenik et al., 2023) are based on variants of the UFM (Mixon et al., 2020). The work in (Tirer et al., 2023) attempts to generalize the UFM by adding a penalty term to the loss that ensures that the features matrix is in the vicinity of a predefined matrix. Yet, the model still lacks an explicit connection between the features and the data. In (Wang et al., 2023), the authors avoid optimizing the features directly but assume that the model is linear and that the data is nearly orthogonal, which are restrictive assumptions. In (Seleznova et al., 2023), the authors claim that having an exact class-wise block structure in the Gram matrix of the empirical NTK on training samples implies NC. Yet, they do not provide reasoning for reaching this collapsed Gram matrix and the analysis is still disconnected from the data.
Here, however, we fully depart from the UFM approach and provide analysis that explicitly depends on the data. Furthermore, unlike most works, our analysis applies to the less studied case where the data is class imbalanced (Fang et al., 2021; Yang et al., 2022; Dang et al., 2023; Thrampoulidis et al., 2022). Our kernel-based analysis utilizes results on NNGP and NTK from (Williams, 1996; Cho & Saul, 2009; Lee et al., 2018; Jacot et al., 2018; Lee et al., 2019). To simplify the analysis, we theoretically analyze kernels associated with shallow fully connected NNs. Focusing on shallow networks is justified by recent works demonstrating monotonic depthwise evolution of NC1 (Tirer & Bruna, 2022; Rangamani et al., 2023; Sukenik et al., 2023; He & Su, 2023; Tirer et al., 2023). Specifically, a data structure that promotes a larger reduction in NC1 for shallow NNs is expected to be more collapsed when using deep NNs. Since NC is related to training, and we show the limitation of NTK to capture it when compared to NNGP, we also utilize the generalization of NNGP that has been proposed in (Seroussi et al., 2023). In this data-aware GP kernel approach, there is an explicit kernel function that depends on the training data. Recently, this richer model has been used to study phase transition behaviors, such as grokking (Rubin et al., 2024), that cannot be captured with the data-independent kernels.
@@ 3 Problem Setup
$\bullet$ Neural Network: Unless stated otherwise, we consider a 2-layer fully connected neural network (2L-FCN) $\psi:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{2}}$ with $l^{th}$ layer width $d_{l},l\in\{1,2\}$, and point-wise activation function $\phi(\cdot):\mathbb{R}\to\mathbb{R}$. Let $\mathbf{W}^{(l)}\in\mathbb{R}^{d_{l}\times d_{l-1}},\mathbf{b}^{(l)}\in\mathbb{R}^{d_{l}}$ denote the weight and bias parameters of the $l^{th}$ layer. At initialization, the entries $W_{ij}^{(l)},b_{i}^{(l)}$ are drawn i.i.d from Gaussian distributions of mean 00 0 and variance $\sigma_{w}^{2}/d_{l-1},\sigma_{b}^{2}$, respectively. For an input $\mathbf{x}\in\mathbb{R}^{d_{0}}$ to the network $\psi(\cdot)$, we denote the $i^{th}$ component of the output vector $\hat{y}_{i}(\mathbf{x})\in\mathbb{R}$ as:
@@
$\displaystyle\hat{\mathbf{Y}}=\begin{bmatrix}\hat{\mathbf{y}}^{1,1}&\cdots&\hat{\mathbf{y}}^{C,n_{C}}\end{bmatrix},\hskip 20.0pt\mathbf{Y}=\begin{bmatrix}\mathbf{y}^{1,1}&\cdots&\mathbf{y}^{C,n_{C}}\end{bmatrix}$ (2)
@@
Note that training classifiers with MSE loss has been shown to be a useful strategy (Hui & Belkin, 2021; Han et al., 2022), and is commonly considered in NC analyses (Han et al., 2022; Tirer & Bruna, 2022).
@@ 4 Within-Class Variability Metric (NC1) for Kernels
In this section, we derive an NC1 metric that depends on the features only through the kernel function. Let $\mathbf{H}\in\mathbb{R}^{N\times d}$ be a matrix encapsulating arbitrary feature vectors associated with samples of the $C$ classes. We define the within-class covariance $\boldsymbol{\Sigma}_{W}(\mathbf{H})$ and between-class covariance $\boldsymbol{\Sigma}_{B}(\mathbf{H})$ matrices of the features as:
@@
Based on these formulations, we define the variability metric $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$, introduced in (Tirer et al., 2023) and used also in (Kothapalli et al., 2023; Wang et al., 2023; Yaras et al., 2023), as:
@@
5 Activation Variability of NNGP and NTK
@@ 5.1 Limiting NNGP Kernel
Under the NN model and initialization stated in Section 3, as the hidden layer width $d_{1}\to\infty$, we can characterize the pre-activation kernel $K^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})$ in terms of the GP limit (Neal, 1995) (commonly referred to as the NNGP limit (Lee et al., 2018)) as follows:
@@
In this limit, the post-activation kernel $Q^{(1)}_{GP}(\cdot,\cdot)$ can have a closed form representation depending on the activation function $\phi(\cdot)$ (Lee et al., 2018; 2019). The expression for Erf activation (Williams, 1996) is:
@@
The formulation for ReLU-based kernel $Q_{GP-ReLU}^{(1)}(\cdot,\cdot)$ (Cho & Saul, 2009) is presented in the Appendix B. Given a kernel function $Q(\cdot,\cdot)$ and samples $\mathbf{X}$, we can formulate the kernel Gram matrix $\mathbf{Q}\in\mathbb{R}^{N\times N}$ as:
@@ 5.2 Limiting NTK
In the infinite width limit, we also analyze the NTK (Jacot et al., 2018) to understand the effect of optimization on the NN’s features in the “lazy regime”. Specifically, a well-known result is that in the infinite width limits (with initialization as per Section 3), the deepest feature mapping of the NN is fixed during gradient descent optimization with small enough learning rate, and is characterized by the NTK. Formally, the recursive relationship between the NTK and NNGP (Jacot et al., 2018; Lee et al., 2019) can be given as follows:
@@
Here, $K_{GP}^{(2)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})$ can be defined using the recursive formulation (Lee et al., 2018; Jacot et al., 2018):
@@
Remark on the limiting NTK.
Denote by $w,\phi$ the parameters and the activation function of an $L$-layer NN $\psi(\cdot)$. The limiting NNGP is defined directly on the product of neurons, $\mathbb{E}_{w}\langle\phi(z_{i}(\mathbf{x})),\phi(z_{i}(\tilde{\mathbf{x}}))\rangle$, and provides kernel expressions for the inner product of features at different layers of the network (via recursion similar to (11)). On the other hand, the limiting NTK is defined as the inner product of the output’s gradients $\mathbb{E}_{w}\langle\nabla_{w}\psi(\mathbf{x}),\nabla_{w}\psi(\tilde{\mathbf{x}})\rangle$. The NTK theory shows that this can model the inner product only of the deepest features (i.e., the output of the penultimate layer) [^1] Note that we denote NNGP $(Q_{GP}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j}))$ and NTK ($\Theta^{(2)}_{NTK}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})$) with different superscripts to be consistent with the literature. Yet both are associated with the output of the single hidden layer. . Yet, NC particularly considers the deepest features.
5.3 On the Limitations of NTK to Exhibit ‘More’ Collapsed Features than NNGP